# ReVISE: Learning to Refine at Test-Time via Intrinsic Self-Verification

### v1

@@ 1 Introduction
Contribution. We propose Refine Via Intrinsic SElf-Verification (ReVISE), a novel and effective self-correction framework for LLM reasoning using self-verification. The core idea of ReVISE is to enable LLMs to assess their reasoning process and refine reasoning trajectories based on self-verification. Specifically, we introduce a special token, which outputs whether to stop the generation or revise the reasoning trajectory. To train the model to utilize this token effectively, we design a two-stage curriculum to simplify the learning of two challenging tasks—self-verification and self-correction—by breaking them into separate training stages. Here, both stages employ preference learning, allowing the model to learn these tasks efficiently without heavy computational overhead. In the first stage, we collect pairs of correct and incorrect reasoning trajectories (i.e., positive and negative samples for preference learning) based on output correctness to develop the model’s self-verification ability. In the second stage, we generate new preference pairs for self-correction by constructing positive samples where a correct reasoning path follows an incorrect one, and negative samples where an incorrect reasoning path follows a correct one.
@@ 2 Related Work
In this section, we provide a comprehensive review of related works, including reasoning, test-time scaling, and self-improvement for LLMs.
@@ 3.3 Verification Confidence-Aware Sampling
GSM8K MATH-500 MATH-500
Methods Maj@1 Maj@3 Maj@5 Maj@1 Maj@3 Maj@5 Maj@1 Maj@3 Maj@5
Few-shot CoT 05.7 06.8 07.2 03.0 02.6 03.2 23.4 22.2 23.2
SFT (Brown et al., 2020a) 22.1 23.5 26.4 10.4 10.6 11.4 27.8 31.0 33.2
RFT (Yuan et al., 2023) 26.2 26.8 28.6 12.6 12.4 12.8 30.8 33.2 35.6
STaR+ (Zelikman et al., 2022) 26.2 27.1 29.9 11.4 13.1 13.4 30.4 31.8 32.8
ReVISE (Ours) 28.1 31.3 32.8 13.4 14.0 14.8 33.6 36.0 37.6
@@ 4 Experiments
Evaluation setup. We mainly report Majority Voting at N (Maj@N) as a sampling-based metric, exceptionally ReVISE used verification confidence-aware majority voting as described in Section 3.3 (unless otherwise specified). We evaluate ReVISE and baselines on GSM8K (Cobbe et al., 2021) and MATH-500 (Hendrycks et al., 2021), a widely used evaluation benchmark subset of MATH.
4.1 Main Results
@@
We first present the main result by comparing the math problem-solving performance with other baselines. Here, we mainly compare ReVISE with various fine-tuning schemes that use a single network and do not use reinforcement learning. Furthermore, we demonstrate the performance of each method with simple test-time scaling methods (i.e., majority voting for baseline methods and using our verification-aware sampling for ReVISE).
@@ Appendix A Experimental Details
  • GSM8K. Used the test split as a benchmark dataset.
@@ Appendix B Additional Results
B.1 Iterative refining sequentially at test time.
(a) Llama-3.2-8B fine-tuned in MATH and evaluated at MATH-500
(b) Llama-3.2-1B-instruct evaluated at GSM8K
Although ReVISE is trained to refine its output in a single, we explore its potential for iterative refinement. Specifically, after generating the second attempt, we append it to the original prompt $x$ and treat it as the first-attempt output. This allows the model to either output $[\mathtt{eos}]$ to terminate the sequence or generate a third attempt following the same process, effectively enabling multiple rounds of refinement. As shown in Figure 8, in MATH-500, ReVISE trained on Llama-3.1-8B’s accuracy consistently improves as the model iteratively refines its responses up to 3 times.
This observation suggests the potential for training a model explicitly designed for multi-round iterative refinement to enhance the iterative refinement even more. By incorporating iterative refinement directly into the training process, the model could learn to self-correct more effectively across multiple rounds. We leave this direction as an exciting avenue for future work.
B.2 ReVISE Results Example
Refinement via intrinsic self-verification example
GSM8K test sample ReVISE trained on Llama-3.2-1B
@@ B.2 ReVISE Results Example
Initial Response The price of carrots increases by 5/100 * \$120 = \$6. So the price of carrots after one year is $120+$6 = \$126. The price of carrots after two years is $126+$6 = \$132. The answer is: 132
Verify $[\mathtt{refine}]$
Refined Response The price of carrots increases by 5/100*\$120 = \$6 per year.The price after one year is \$120 + \$6 = \$126 The price after two years is \$126 + \$6 = \$132. The price after three years is \$132 + \$6 = \$138. The answer is: 138
@@
MATH test sample ReVISE trained on Llama-3.2-1B
@@
Initial Response We can set up a system of equations to solve this problem. Let $t$ be the weight of one treek, $s$ be the weight of one squig, and $g$ be the weight of one goolee. From the given information, we can write the following equations: $10t=3s+g$ $2t+g=s$ We can solve this system of equations to find the weight of one treek in terms of the weight of one squig. Multiplying the second equation by 3, we get: $6t+3g=3s$ Subtracting the first equation from this equation, we get: $6t+3g-10t=3s-(3s+g)$ Simplifying, we get: $-4t+3g=-g$ Adding $4t$ to both sides, we get: $3g=4t$ Dividing both sides by 3, we get: $g=\frac{4}{3}t$ Substituting this into the second equation, we get: $2t+\frac{4}{3}t=s$ Multiplying both sides by 3, we get: $6t+4t=3s$ Combining like terms, we get: $10t=3s$ Since we know that $10t=3s+g$, we can substitute $g=\frac{4}{3}t$ into this equation: $10t=3s+\frac{4}{3}t$ Multiplying both sides by 3, we get: $30t=9s+4t$ Subtracting $4t$ from both sides, we get: $26t=9s$ Dividing both sides by 9, we get: $\frac{26}{9}t=s$ So, the weight of one treek is equal to $\frac{9}{26}$ times the weight of one squig. Therefore, the combined weight of how many treeks equals the weight of one squig is $\frac{9}{26}$. The answer is: $\boxed{\frac{9}{26}}$
Verify $[\mathtt{refine}]$
@@
Stop via intrinsic self-verification example
GSM8K test sample ReVISE trained on Llama-3.2-1B
@@
Initial Response Jen earns 6 x \$1.5 = \$9 for the yams. She earns 10 x \$2 = \$20 for the sweet potatoes. She earns 4 x \$1.25 = \$5 for the carrots. Therefore, Jen earns \$9 + \$20 + \$5 = \$34 in all. The answer is: 34
Verify $[\mathtt{eos}]$
@@
MATH test sample ReVISE trained on Llama-3.2-1B
@@
Initial Response $(1+2i)6-3i=6+12i-3i=6+9i=\boxed{6+9i}$. The answer is: $\boxed{6+9i}$.
Verify $[\mathtt{eos}]$
@@ Appendix C The few-shot prompt used for evaluation
Few-shot prompt for GSM8K:
Given the following problem, reason and give a final answer to the problem. Problem: There are 15 trees in the grove. Grove workers will plant trees in the grove today. After they are done, there will be 21 trees. How many trees did the grove workers plant today? Your response should end with ”The final answer is [answer]” where [answer] is the response to the problem. There are 15 trees originally. Then there were 21 trees after some more were planted. So there must have been 21 - 15 = 6. The final answer is 6 Given the following problem, reason and give a final answer to the problem. Problem: If there are 3 cars in the parking lot and 2 more cars arrive, how many cars are in the parking lot? Your response should end with ”The final answer is [answer]” where [answer] is the response to the problem. There are originally 3 cars. 2 more cars arrive. 3 + 2 = 5. The final answer is 5 Given the following problem, reason and give a final answer to the problem. Problem: Leah had 32 chocolates and her sister had 42. If they ate 35, how many pieces do they have left in total? Your response should end with ”The final answer is [answer]” where [answer] is the response to the problem. Originally, Leah had 32 chocolates. Her sister had 42. So in total they had 32 + 42 = 74. After eating 35, they had 74 - 35 = 39. The final answer is 39 Given the following problem, reason and give a final answer to the problem. Problem: Jason had 20 lollipops. He gave Denny some lollipops. Now Jason has 12 lollipops. How many lollipops did Jason give to Denny? Your response should end with ”The final answer is [answer]” where [answer] is the response to the problem. Jason started with 20 lollipops. Then he had 12 after giving some to Denny. So he gave Denny 20 - 12 = 8. The final answer is 8 Given the following problem, reason and give a final answer to the problem. Problem: Shawn has five toys. For Christmas, he got two toys each from his mom and dad. How many toys does he have now? Your response should end with ”The final answer is [answer]” where [answer] is the response to the problem. Shawn started with 5 toys. If he got 2 toys each from his mom and dad, then that is 4 more toys. 5 + 4 = 9. The final answer is 9 Given the following problem, reason and give a final answer to the problem. Problem: There were nine computers in the server room. Five more computers were installed each day, from monday to thursday. How many computers are now in the server room? Your response should end with ”The final answer is [answer]” where [answer] is the response to the problem. There were originally 9 computers. For each of 4 days, 5 more computers were added. So 5 * 4 = 20 computers were added. 9 + 20 is 29. The final answer is 29 Given the following problem, reason and give a final answer to the problem. Problem: Michael had 58 golf balls. On tuesday, he lost 23 golf balls. On wednesday, he lost 2 more. How many golf balls did he have at the end of wednesday? Your response should end with ”The final answer is [answer]” where [answer] is the response to the problem. Michael started with 58 golf balls. After losing 23 on tuesday, he had 58 - 23 = 35. After losing 2 more, he had 35 - 2 = 33 golf balls. The final answer is 33 Given the following problem, reason and give a final answer to the problem. Problem: Olivia has \$23. She bought five bagels for \$3 each. How much money does she have left? Your response should end with ”The final answer is [answer]” where [answer] is the response to the problem. Olivia had 23 dollars. 5 bagels for 3 dollars each will be 5 x 3 = 15 dollars. So she has 23 - 15 dollars left. 23 - 15 is 8. The final answer is 8 Given the following problem, reason and give a final answer to the problem. Problem: Janet’s ducks lay 16 eggs per day. She eats three for breakfast every morning and bakes muffins for her friends every day with four. She sells the remainder at the farmers’ market daily for \$2 per fresh duck egg. How much in dollars does she make every day at the farmers’ market? Your response should end with ”The final answer is [answer]” where [answer] is the response to the problem.
Few-shot prompt for MATH-500:

### v2

@@ 1 Introduction
Contribution. We propose Refine Via Intrinsic SElf-Verification (ReVISE) [^1] Code available at: github.com/seunghyukoh/revise, a novel and effective self-correction framework for LLM reasoning using self-verification. The core idea of ReVISE is to enable LLMs to assess their reasoning process and refine reasoning trajectories based on self-verification. Specifically, we introduce a special token, which outputs whether to stop the generation or revise the reasoning trajectory. To train the model to utilize this token effectively, we design a two-stage curriculum to simplify the learning of two challenging tasks—self-verification and self-correction—by breaking them into separate training stages. Here, both stages employ preference learning, allowing the model to learn these tasks efficiently without heavy computational overhead. In the first stage, we collect pairs of correct and incorrect reasoning trajectories (i.e., positive and negative samples for preference learning) based on output correctness to develop the model’s self-verification ability. In the second stage, we generate new preference pairs for self-correction by constructing positive samples where a correct reasoning path follows an incorrect one, and negative samples where an incorrect reasoning path follows a correct one.
@@ 3.3 Verification Confidence-Aware Sampling
GSM8K MATH-500 GSM8K MATH-500
Methods Maj@1 Maj@5 Maj@1 Maj@5 Maj@1 Maj@5 Maj@1 Maj@5
Few-shot CoT 05.7 07.2 03.0 03.2 56.7 58.3 23.4 23.2
SFT (Brown et al., 2020a) 22.1 26.4 10.4 11.4 58.2 64.8 27.8 33.2
RFT (Yuan et al., 2023) 26.2 28.6 12.6 12.8 58.9 65.3 30.8 35.6
STaR+ (Zelikman et al., 2022) 26.2 29.9 11.4 13.4 59.2 64.9 30.4 32.8
ReVISE (Ours) 28.1 32.8 13.4 14.8 61.6 69.2 33.6 37.6
@@ 4 Experiments
Evaluation setup. We mainly report Majority Voting at K (Maj@K) as a sampling-based metric, exceptionally ReVISE used verification confidence-aware majority voting as described in Section 3.3 (unless otherwise specified). We evaluate ReVISE and baselines on GSM8K (Cobbe et al., 2021) and MATH-500 (Hendrycks et al., 2021), a widely used evaluation benchmark subset of MATH.
@@
4.1 Main Results
We first present the main result by comparing the math problem-solving performance with other baselines. Here, we mainly compare ReVISE with various fine-tuning schemes that use a single network and do not use reinforcement learning. Furthermore, we demonstrate the performance of each method with simple test-time scaling methods (i.e., majority voting for baseline methods and using our verification-aware sampling for ReVISE). Also, we verify that ReVISE effectively enhances reasoning in the coding domain (i.e., MBPP (Austin et al., 2021)).
@@ 4.1 Main Results
Table 2: Results on the MBPP (Austin et al., 2021) benchmark for ReVISE and baselines trained on Llama-3.2-1B.
Method Pass@1
Few-shot CoT 24.5
SFT 30.0
RFT 29.6
STaR+ 30.7
ReVISE (Ours) 33.1
As shown in Table 2, we further investigate the performance of ReVISE in the coding benchmark MBPP (Austin et al., 2021). Specifically, ReVISE surpasses all baseline approaches, achieving a Pass@1 score of 33.1%, notably outperforming strong baselines such as SFT (30.0%) and STaR+ (30.7%). These results highlight ReVISE’s effectiveness beyond mathematical reasoning, extending its superior refinement capabilities to code-generation tasks as well. The consistent performance improvement across diverse benchmarks underscores the generalizability and robustness of the intrinsic refinement strategy employed by ReVISE.
@@ 4.3 Additional Analysis and Ablation
(a) Llama-3.2-8B fine-tuned in MATH and evaluated at MATH-500
(b) Llama-3.2-1B-instruct evaluated at GSM8K
@@
Table 4: Accuracy (%) on GSM8K under transfer domain generalization. All models are trained on MATH and evaluated on GSM8K. Trained on Llama-3.2-1B and Llama-3.1-8B models.
Model Methods Accuracy (%)
Llama-3.2-1B SFT 7.3
RFT 8.2
STaR+ 8.0
ReVISE (Ours) 8.8
Llama-3.1-8B SFT 60.3
RFT 60.3
STaR+ 58.7
ReVISE (Ours) 61.5
Generalization under transfer dataset domain. We demonstrate the generalization ablity of ReVISE in a transfer domain setting (see Table 4). Specifically, we train ReVISE on the MATH domain and test on the GSM8K domain, using both Llama-3.2-1B and Llama-3.1-8B models. As shown in Table 4, both model sizes significantly outperform other baselines in this out-of-distribution evaluation. For example, ReVISE achieves an accuracy of 8.8% using Llama-3.2-1B, and 61.5% using Llama-3.1-8 B. These results demonstrate that our method possesses strong domain transferability, maintaining its advantage over baselines even when evaluated on a different dataset.
Iterative refining sequentially at test time. Although ReVISE is trained to refine its output in a single pass, we explore its potential for iterative refinement. Specifically, after generating the second attempt, we append it to the original prompt $x$ and treat it as the first-attempt output. This allows the model to either output $[\mathtt{eos}]$ to terminate the sequence or generate a third attempt following the same process, effectively enabling multiple rounds of refinement. As shown in Figure 8, in MATH-500, ReVISE trained on Llama-3.1-8B’s accuracy consistently improves as the model iteratively refines its responses up to 3 times.
This observation suggests the potential for training a model explicitly designed for sequential iterative refinement to enhance the iterative refinement even more. By incorporating iterative refinement directly into the training process, the model could learn to self-correct more effectively across multiple rounds. We leave this direction as an exciting avenue for future work.
@@ 5 Conclusion
Acknowledgements
This work was supported by Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II190075 Artificial Intelligence Graduate School Program(KAIST); No. RS-2024-00509279, Global AI Frontier Lab) and NIPA(National IT Industry Promotion Agency), through the Ministry of Science and ICT (Hyperscale AI flagship project).
@@ Appendix A Experimental Details
  • MBPP. MBPP (Austin et al., 2021) is a collection of crowd-sourced Python programming problems. Each instance consists of a natural language task description, a reference solution, and three test cases written in Python. Since ReVISE requires intermediate reasoning steps not provided in the original dataset, we generated them by applying Chain-of-Thought prompting to GPT-4o. For each problem, we collected 16 valid reasoning paths along with corresponding code solutions that successfully pass all test cases.
@@
  • GSM8K, MBPP. Used the test split as a benchmark dataset.
@@ Appendix B Additional Results
B.1 Comparison with DPO
Table 5: Comparison between DPO and ReVISE. We report accuracy (%) on GSM8K and MATH-500. The models are trained on Llama-3.2-1B. The bold indicates the best result within the group.
Method GSM8K MATH-500
DPO 22.6 10.8
ReVISE (Ours) 28.1 13.4
To further analyze the effectiveness of ReVISE’s training framework, we include a comparison with a reinforcement learning(RL)-based baseline using Direct Preference Optimization (DPO). We trained the DPO model from the same supervised fine-tuned checkpoint, which was used in training ReVISE. We construct a preference pair dataset where ground truth answers are preferred over incorrect responses. As shown in Table 5, ReVISE even outperforms the DPO trained baseline, indicating 22.7% in GSM8K and 10.8% in MATH-500.
B.2 Iterative Training and Comparison with Self-Correction Work.
Table 6: Comparison between ReVISE and other baselines (i.e., Zero-shot CoT, SFT, RFT, and SCoRe (Kumar et al., 2024)). We report accuracy (%) on MATH-500. The models are trained on Gemma-2-2B. The bold indicates the best result within the group.
Method Accuracy (%) Training Efficiency
Zero-shot CoT 16.8 -
SFT 17.6 -
RFT 18.6 -
SCoRe 23.0 x1
ReVISE (Ours) 23.2 x30
+ iter1 (Ours) 24.2 x20
+ iter2 (Ours) 25.8 x15
We include a comparison with the self-correction baseline, SCoRe (Kumar et al., 2024). While SCoRe relies on costly online reinforcement learning and requires extensive reasoning path generation (resulting in approximately 1.5 million generations for 3,000 steps with a batch size of 512), ReVISE is significantly more efficient. Specifically, ReVISE constructs preference pairs by generating a single reasoning path per sample, totaling only 50,000 generations for the entire dataset. This corresponds to a 30× reduction in training cost compared to SCoRe. Despite this large efficiency gap, ReVISE achieves higher accuracy than SCoRe on the MATH-500 benchmark using the same Gemma2-2B model, as shown in Table 6. Furthermore, ReVISE ’s performance improves further with repeated training cycles. At each cycle, we re-sample reasoning path pairs with the current model and iteratively apply preference optimization to refine it further. This strategy leads to continual accuracy gains while maintaining substantial efficiency benefits. For instance, after two additional iterations, ReVISE achieves 25.8% accuracy, which further increases its performance margin over SCoRe while incurring a training cost that is 15 times smaller.
These results demonstrate that ReVISE is not only more practical and scalable for self-correction tasks, but also able to leverage iterative refinement to reach even higher performance. All comparisons utilize SCoRe’s results from the original paper, as there is a lack of open-source code.
B.3 Quantify the Verifying Performance.
Table 7: Comparison of AUROC (%) of verifying correctness between ReVISE and V-STaR (Hosseini et al., 2024) verifier. We report AUROC (%) on GSM8K. The models are trained on Llama-3.2-1B.
Method AUROC (%)
V-STaR verifier 69.5
ReVISE (Ours) 76.0
We further evaluate the quality of the verifying signal produced by ReVISE using the area under the receiver operating characteristic curve (AUROC) and compare with V-STaR (Hosseini et al., 2024) verifier, while we reported ReVISE’s calibration performance using AUROC in Section 4 and Figure 3(b). As shown in Table 7, ReVISE achieves an AUROC of 76.0%, outperforming the V-STaR verifier, which achieves 69.5%, even though V-STaR uses a separately trained verifier. These experimental results demonstrate the effectiveness of ReVISE’s intrinsic verifier, resulting in improved performance and enhanced test-time scalability.
B.4 Extended Test-time Scaling Behavior Experiment.
Table 8: Test-time scaling results (Maj@K) on GSM8K with Llama-3.2-1B. We evaluate accuracy (%) as the number of sampled generations K increases from 2 to 64. We compared ReVISE and other baselines (i.e., SFT, RFT, and STaR+).
Maj@2 Maj@4 Maj@8 Maj@16 Maj@32 Maj@64
SFT 20.5% $\pm$ 0.5 24.5% $\pm$ 0.6 28.2% $\pm$ 0.6 30.0% $\pm$ 0.4 31.8% $\pm$ 0.1 32.1% $\pm$ 0.4
RFT 24.6% $\pm$ 0.3 27.5% $\pm$ 0.3 29.8% $\pm$ 0.4 30.9% $\pm$ 0.5 31.3% $\pm$ 0.3 33.2% $\pm$ 0.2
STaR+ 24.0% $\pm$ 1.1 27.1% $\pm$ 0.6 29.3% $\pm$ 0.6 30.4% $\pm$ 0.6 31.1% $\pm$ 0.5 31.6% $\pm$ 0.4
ReVISE (Ours) 28.3% $\pm$ 0.7 32.5% $\pm$ 0.9 34.9% $\pm$ 0.5 36.2% $\pm$ 0.4 37.2% $\pm$ 0.5 37.7% $\pm$ 0.5
We extended our experiments on test-time scaling behavior of ReVISE and other baselines (i.e., SFT, RFT, and STaR+) on GSM8K using the Llama-3.2-1B models. Specifically, for the baselines, we measure accuracy by majority voting over K sampled generations, where K ranges from 2 to 64 (Maj@K). For ReVISE, we used our own method, confidence-based weighted majority voting, for the same generation ranges. As shown in Table 8, ReVISE consistently outperforms every baseline across every sample size. Notably, ReVISE reaches 37.7% accuracy at Maj@64, representing a substantial improvement over prior methods. These results demonstrate the effectiveness of ReVISE in leveraging increased generation diversity for robust test-time self-verification.
B.5 ReVISE Results Example
B.5.1 Refinement via intrinsic self-verification example
Table 9: An example of ReVISE’s refinement on a GSM8K question
@@ B.5.1 Refinement via intrinsic self-verification example
[…]
@@
Table 10: An example of ReVISE’s refinement on a MATH question
@@
B.5.2 Stop via intrinsic self-verification example
Table 11: An example where ReVISE chooses not to refine on a GSM8K question
@@ B.5.2 Stop via intrinsic self-verification example
[…]
@@
Table 12: An example where ReVISE chooses not to refine on a MATH question
@@
Response $(1+2i)6-3i=6+12i-3i=6+9i$. The answer is: $\boxed{6+9i}$. $[\mathtt{eos}]$
@@ Appendix C The few-shot prompt used for evaluation
Tables 13-15 present the few-shot templates used for evaluation. We filled in the {placeholders} using using the questions, answers, and test cases (for MBPP only) from Tables 16-C.
Table 13: The few-shot template used for GSM8K
Template for GSM8K
Given the following problem, reason and give a final answer to the problem. Problem: {question} Your response should end with ”The final answer is [answer]” where [answer] is the response to the problem. {answer}.
Table 14: The few-shot template used for MATH
Template for MATH-500
Problem: {question} Answer: {answer}.
Table 15: The few-shot template used for MBPP
Template for MBPP
You are given a programming problem. Let’s reason step by step before writing the code. Think through the problem carefully, explain your reasoning clearly, and then at the very end, provide your final code.
Here is your task: {question}
Your code should pass these tests, and do not include the following test code in your Python code:
{test cases}
{answer}
Table 16: The 8 few-shot examples used for evaluation on GSM8K
Question Answer
There are 15 trees in the grove. Grove workers will plant trees in the grove today. After they are done, there will be 21 trees. How many trees did the grove workers plant today? There are 15 trees originally. Then there were 21 trees after some more were planted. So there must have been 21 - 15 = 6. The final answer is 6
If there are 3 cars in the parking lot and 2 more cars arrive, how many cars are in the parking lot? There are originally 3 cars. 2 more cars arrive. 3 + 2 = 5. The final answer is 5
Leah had 32 chocolates and her sister had 42. If they ate 35, how many pieces do they have left in total? Originally, Leah had 32 chocolates. Her sister had 42. So in total they had 32 + 42 = 74. After eating 35, they had 74 - 35 = 39. The final answer is 39
Jason had 20 lollipops. He gave Denny some lollipops. Now Jason has 12 lollipops. How many lollipops did Jason give to Denny? Jason started with 20 lollipops. Then he had 12 after giving some to Denny. So he gave Denny 20 - 12 = 8. The final answer is 8
Shawn has five toys. For Christmas, he got two toys each from his mom and dad. How many toys does he have now? Shawn started with 5 toys. If he got 2 toys each from his mom and dad, then that is 4 more toys. 5 + 4 = 9. The final answer is 9
There were nine computers in the server room. Five more computers were installed each day, from mondax“y to thursday. How many computers are now in the server room? There were originally 9 computers. For each of 4 dayxx‘s, 5 more computers were added. So 5 * 4 = 20 computers were added. 9 + 20 is 29. The final answer is 29
Michael had 58 golf balls. On tuesday, he lost 23 golf balls. On Wednesday, he lost 2 more. How many golf balls did he have at the end of wednesday? Michael started with 58 golf balls. After losing 23 on tuesday, he had 58 - 23 = 35. After losing 2 more, he had 35 - 2 = 33 golf balls. The final answer is 33
Olivia has \$23. She bought five bagels for \$3 each. How much money does she have left? Olivia had 23 dollars. 5 bagels for 3 dollars each will be 5 x 3 = 15 dollars. So she has 23 - 15 dollars left. 23 - 15 is 8. The final answer is 8
Table 17: The 4 few-shot examples used for evaluation on MATH
Question Answer
Find the domain of the expression $\frac{\sqrt{x-2}}{\sqrt{5-x}}.$ The expressions inside each square root must be non-negative. Therefore, $x-2\geq 0$, so $x\geq 2$, and $5-x\geq 0$, so $x\leq 5$. Also, the denominator cannot be equal to zero, so $05-x>$ 0, which gives $x<5$. Therefore, the domain of the expression is $\boxed{[2,5)}$. The answer is $\boxed{[2,5)}$.
@@
[…]
If the system of equations $6x-4y=a,6y-9x=b,$ has a solution $(x,y)$ where $x$ and $y$ are both nonzero, find $\frac{a}{b}$, assuming $b$ is nonzero. If we multiply the first equation by $-\frac{3}{2}$, we obtain $6y-9x=-\frac{3}{2}a.$ Since we also know that $6y-9x=b$, we have $-\frac{3}{2}a=b\Rightarrow\frac{a}{b}=\boxed{-\frac{2}{3}}.$ The answer is $\boxed{-\frac{2}{3}}$.
Table 18: The 3 few-shot examples used for evaluation on MBPP
Question Test Cases Answer
Write a function to find the similar elements from the given two tuple lists. assert similar_elements((3, 4, 5, 6),(5, 7, 4, 10)) == (4, 5) assert similar_elements((1, 2, 3, 4),(5, 4, 3, 7)) == (3, 4) assert similar_elements((11, 12, 14, 13),(17, 15, 14, 13)) == (13, 14) Step 1: Analyze the Problem We are given two tuples of integers and asked to find the ”similar elements” between them. This implies we need to identify the elements that are common to both tuples. The result should be returned as a tuple, and the order of elements in the result should be sorted in ascending order, based on the test cases. Step 2: Plan the Solution 1. Input: Two tuples, e.g., ‘(3, 4, 5, 6)‘ and ‘(5, 7, 4, 10)‘. 2. Goal: Find common elements between the two tuples. 3. Output: A tuple of the common elements in sorted order. Steps: Convert both tuples into sets. This allows us to perform set intersection efficiently. Use the ‘&‘ operator or ‘.intersection()‘ method to get common elements. Convert the resulting set to a sorted list. Convert that sorted list back into a tuple. Return the result. Step 3: Consider Edge Cases If there are no common elements, the result should be an empty tuple: ‘()‘. The order of elements in the result must be sorted ascendingly. Now that the reasoning is clear, let’s move on to the code. The answer is: {minted}python def similar_elements(t1, t2): return tuple(sorted(set(t1) set(t2)))
Write a function to identify non-prime numbers. assert is_not_prime(2) == False assert is_not_prime(10) == True assert is_not_prime(35) == True Step 1: Analyze the Problem We are asked to write a function ‘is_not_prime(n)‘ that determines whether a number is not a prime. The function should return: ‘True‘ if the number is not a prime. ‘False‘ if the number is a prime. Step 2: Understand the Requirements A prime number is a number greater than 1 that has no divisors other than 1 and itself. So, the function should: Return ‘True‘ if ‘n‘ is less than 2 (since 0 and 1 are not prime). Return ‘True‘ if ‘n‘ has any divisor other than 1 and itself. Return ‘False‘ otherwise. Step 3: Plan the Solution 1. Handle edge cases: If ‘n ¡= 1‘, return ‘True‘ because these are not prime. 2. Loop from ‘2‘ to ‘sqrt(n)‘: - If any number in this range divides ‘n‘, it is not prime → return ‘True‘. 3. If the loop completes without finding a divisor, return ‘False‘. Step 4: Write and Test the Code The answer is: {minted}python import math def is_not_prime(n): if n ¡= 1: return True for i in range(2, int(math.sqrt(n)) + 1): if n return True return False
Write a function to find the largest integers from a given list of numbers using heap queue algorithm. assert heap_queue_largest( [25, 35, 22, 85, 14, 65, 75, 22, 58],3)==[85, 75, 65] assert heap_queue_largest( [25, 35, 22, 85, 14, 65, 75, 22, 58],2)==[85, 75] assert heap_queue_largest( [25, 35, 22, 85, 14, 65, 75, 22, 58],5)==[85, 75, 65, 58, 35] Step 1: Analyze the Problem We are given a list of integers and a number ‘n‘, and we need to find the ‘n‘ largest integers from that list using a heap queue algorithm. This means: We should use a heap, which is a special kind of binary tree used primarily for priority queues. Python provides a built-in module called ‘heapq‘ which supports heap operations. Key point: Although ‘heapq‘ is a min-heap by default, it provides utility functions to get the ‘n‘ largest elements efficiently: ‘heapq.nlargest(n, iterable)‘. Step 2: Plan the Solution 1. Import the ‘heapq‘ module. 2. Use ‘heapq.nlargest(n, iterable)‘ to get the largest ‘n‘ elements from the list. 3. Return that list. Step 3: Write the Code This will be a single function ‘heap_queue_largest(nums, n)‘ that uses ‘heapq.nlargest‘. The answer is: {minted}python import heapq def heap_queue_largest(nums, n): return heapq.nlargest(n, nums)